OpenStudy (wormynut):

HELP QUICK! WILL GIVE MEDALS! Give the consecutive odd, consecutive even, consecutve positive and consecutive negative integers in terms of x.

OpenStudy (wormynut):

yeah....for example: consecutive positive would be x and x+1

OpenStudy (wormynut):

right?

OpenStudy (kittiwitti1):

@skullpatrol

OpenStudy (wormynut):

lol

OpenStudy (wormynut):

I really need the answer quickly!

OpenStudy (wormynut):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

im not familiar with these terms

OpenStudy (kittiwitti1):

@inkyvoyd @jdoe0001

OpenStudy (wormynut):

thanks for trying to help me :) @kittiwitti1

OpenStudy (kittiwitti1):

No problem ☺

OpenStudy (wormynut):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

Let the odd number be x. Then consecutive odd numbef will be x+2

OpenStudy (wormynut):

ok

OpenStudy (jiteshmeghwal9):

similarly even number also have a gap of 2. if an even number is x. then consecutive even number is x+2

OpenStudy (wormynut):

ohhh yeah!

OpenStudy (wormynut):

wait, wont odd be x+2? Let x=1, then 1+2=3?

OpenStudy (jiteshmeghwal9):

yes that's what i wrote

OpenStudy (wormynut):

oh yeah...didn't read properly!

OpenStudy (jiteshmeghwal9):

np :)

OpenStudy (wormynut):

wait, what about positive and negative? @jiteshmeghwal9

OpenStudy (skullpatrol):

What about them?

OpenStudy (wormynut):

are there any equations for that?

OpenStudy (skullpatrol):

Yes, if n is a positive integer then -n is a negative integer.

OpenStudy (wormynut):

oh so what do I write if they ask for even negative integer?

OpenStudy (skullpatrol):

What is your definition of an even integer?

OpenStudy (wormynut):

ummm....divisible by 2?

OpenStudy (jiteshmeghwal9):

consecutive integers are always of the form x+1 take example x=1 x+1=1+1=2 x=-2 x+1=-2+1=-1

OpenStudy (wormynut):

oh okaay THANK YOU LL VERY MUCH!!!!

OpenStudy (jiteshmeghwal9):

yw :)